OpenStudy (anonymous):

Find the equation of the parabola that passes through each set of points. (2,0), (3,-2), (1,-2)

OpenStudy (anonymous):

I was able to get it to the point where I substituted in each point into y=ax^2+bx+c but wasn't able to isolate one variable in order to solve for the rest

hero (hero):

There's a couple different ways this can be done

OpenStudy (anonymous):

Normally they have at least one point that has x as being 0 that way c can be solved for and therefor the rest but that doesn't happen in this case and i am unsure of how to do it

OpenStudy (anonymous):

any ideas?

OpenStudy (dumbcow):

well you end up with system of 3 equations: 4a+2b+c = 0 9a +3b +c = -2 a + b +c = -2 solve using matrices or by hand(elimination)

OpenStudy (anonymous):

how would i do that like multiply the 3rd one by -2 and add it to the 1st one? then do itagain but with a different combination and multiplying by a different number?

hero (hero):

There's an easier way

hero (hero):

The parabola is symmetric so the points (3,-2) and (1,-2) have a midpoint

hero (hero):

you can get the x coordinate of the vertex from that

OpenStudy (dumbcow):

@mskyeg , yes thats the idea Hero is giving easier way, however depending on the points you have to do it the long way

OpenStudy (anonymous):

thanks so much:)

OpenStudy (anonymous):

@Hero thanks for showing me the easier way but I haven't learned that yet and due to the fact I have to show my teacher my work I think she would rather me do it the way she taught but thanks for teaching it to me it'll make future work much easier:)

hero (hero):

I'm curious to know what classroom method you're being taught. I believe the shortcut way was intentional.

OpenStudy (dumbcow):

hmm i agree...the given points are very neatly given (2,0) is vertex with 2 other points equidistant away

OpenStudy (dumbcow):

normally if you are just given 3 random points on parabola then you have to set up system of 3 equations. in this case use vertex form equation y = a(x-h)^2 +k plug in vertex y = a(x-2)^2 plug in point (3,-2) and solve for a

OpenStudy (anonymous):

why (3,-2) what about the other points

OpenStudy (dumbcow):

doesn't matter, you could use other point

OpenStudy (anonymous):

how do i find b?

OpenStudy (anonymous):

and c?

hero (hero):

Use the formula he gave you y = a(x-h)^2 + k

OpenStudy (anonymous):

i did that i only got a what about b and c

hero (hero):

Bro, all you need is y = a(x - h)^2 + k Forget b and c

OpenStudy (anonymous):

it needs to be in the form of ax^2 +bx +c

hero (hero):

a(x-h)^2 + k = ax^2 + bx + c

hero (hero):

They're equal, so just expand whatever you got for a(x-h)^2 + k

OpenStudy (anonymous):

what do i plug in for b andthere is only one other number (-2)

hero (hero):

you have to expand y = a(x-h)^2 + k

hero (hero):

I'll just show you since you have no clue whatsoever.

hero (hero):

We know (2,0) is the vertex, so I just insert it along with point (3,-2) -2 = a(3-2)^2 + 0 -2 = a(3-2)^2 + 0 -2 = a(1)^2 -2 = a(1) -2 = a So y = -2(x-2)^2 + 0 y = -2(x-2)(x-2) y = -2(x(x-2)-2(x-2)) y = -2(x^2 - 2x - 2x - 4) y = -2(x^2 - 4x - 4) y = -2x^2 - 8x - 8

OpenStudy (anonymous):

why is (2,0) the vertex how is that found

hero (hero):

It was given

OpenStudy (anonymous):

it simply said that the parabola passes through those three points it never says which one them if any is the vertex.

hero (hero):

You didn't pay attention to anything @dumbcow said did you?

OpenStudy (anonymous):

got it thanks

hero (hero):

We would not be able to solve this without finding the vertex. It's a key step in solving unless you're going to do systems of equations method. Ugh